Suppose molar specific heat C_V of iron is given by C_V=AT^5, where A is some constant. For iron, A=5.17×〖10〗^(-3) J⁄(mol .K^4 ). Calculate the entropy change for 2.00 mol of iron when its temperature increases from 15 K to 25 K.
1
Expert's answer
2020-08-18T12:26:18-0400

By the definition, "C_V=\frac{dQ}{dT}", where Q is heat, transferred to the 1 mole of material, T - its temperature. Hence, "dQ=C_VdT". The enthropy change "dS=\frac{dQ}{T}". Thus, for n moles: "dS=\frac{n\cdot C_VdT}{T}". If "C_V=AT^5":

"dS=\frac{n\cdot AT^5dT}{T}=n\cdot AT^4dT \Rightarrow ΔS=S_2-S_1=\frac{n\cdot A }{5}\cdot (T^5_2-T^5_1)="

"=\frac {2\cdot 5.17 \cdot 10^{-3}}{5} \cdot (9.77 \cdot 10^{6}-0.76\cdot 10^{6})=18.6 \cdot 10^3 \frac {J}{K}"